OpenStudy (anonymous):

A new stethoscope costs $95. You are able to purchase the stethoscope for $68.40 with your student discount. What was your student discount rate?

OpenStudy (anonymous):

discount= discount rate x list price =$95-$68.40 =26.6 26.6=x(95) 26.6/95=95x/95 x=0.28 x=28%